Find the median of the set of data. (Round your answer to one decimal place.) 230, 220, 212, 212, 212, 212, 220, 215, 215, 215, 220, 220 QUESTION 29 Find the variance of the set of data. (Round your answer to one decimal place.) 230, 220, 212, 212, 212, 212, 220, 215, 215, 215, 220, 220 QUESTION 30 Jackies sisters weigh 114 lb, 140 lb, 120 lb, and 129 lb. The average female in her city weighs 139.1 lb. How much does Jackie weigh if she and her sisters have an average weight of 139.1 lb? (Round the answer to the nearest tenth) Find the range for the given data. (Round to the nearest dollar.) Fred, a local mechanic, gathered the following data regarding the price, in dollars, of an oil and filter change at twelve competing service stations: 32.95, 24.95, 26.95,28.95, 18.95, 28.95, 30.95,22.95, 24.95, 26.95, 29.95, 28.95 QUESTION 25 rvey, students showed these preferences for 12%/ 18% How many degrees are in the central angle for the written sector? (Round answer to the nearest ten
Answer #1

For Question 28 and 29, we can get the values directly using R software using the commands as shown below:

x- c (230, 220,212,212,212,212,220,215,215,215,220, 220) > median (x) [1] 215 > var (x) I1] 28.81061

Q28 : Thus, the median = 215.0 (rounded upto 1 decimal place)

Q29 : variance = 28.8 (rounded upto 1 decimal place)

Q30.

Average weight of Jackie and his four sisters = 139.1 lb

Total weight of Jackie and his four sisters = 5*139.1 lb = 695.5 lb

Weight of Jackie

= Total weight - (Sum of weights of 4 sisters)

= 695.5 - (114+140+120+129) lb

= 192.5 lb

Q24,

Range for the mechanic data

= Maximum value - Minimum Value

= $32.95 -$18.95

=$14 (rounded to nearest dollar)

Q25.

Central Angle for the written sector

= Percentage corresponding to written sector * 360 degrees

=(9/100)*360 degrees

=32.4 degrees (rounded to nearest tenths)